Please help me find the eqation and be able to slove for x.
how many grams of a 50% silver alloy be mixed with a 75% silver alloy to obtain 40 grams of a 60% silver alloy?
I think it's. ( .5x+.75= ( 40*.6)) but I'm not sure

Answer by Alan3354(69443)

Do it like this one. Only the numbers are different.
-----------
Amanda wants to make 8 gal. of a 20% saline solution by mixing together a 56% saline solution and a 8% saline solution. How much of each solution must she use?
================
e = amount of 8%
f = amount of 56%
---
e+ f = 8 (total solution)
8e + 56f = 20*8 (total saline)
---
e+ f = 8
e + 7f = 20
------------ Subtract
-6f = -12
f = 2
e = 6